# Thinking Forward and Backward: Multi-Objective Reinforcement Learning for Retrieval-Augmented Reasoning

### v1

@@ 2 Preliminaries
Search-R1. Recent advances, such as Search-R1, extend RAG to support multi-step reasoning with interleaved retrieval. In this paradigm, given a question $Q$, the LLM generates a reasoning trajectory $T=\{T_{1},T_{2},\ldots,T_{n}\}$. At each step $i$, the LLM (i) first generates a reasoning step $T_{i}$ based on the current information; (ii) then, it issues a search query and retrieves relevant documents; and (iii) finally, it judges whether to move on to the next reasoning step $T_{i+1}$ or generate the final answer $A$ based on the current content. The process alternates between reasoning and search.
@@ 3.2 Bidirectional information quantification
    Step-to-answer distance $d_{\text{T-A}}(T_{i})$ quantifies how much the current step $T_{i}$ contributes toward the final answer $A$, indicating its solution progress; and
@@ 3.3 Multi-objective optimization with RL
represent the rewards derived from the step-to-answer and step-to-question distances at step $i$.
@@
$J_{\text{GRPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot|x;\mathcal{R})}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}$ (8)
@@
$-\beta D_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}]\Bigg],$

### v3

@@ 2 Preliminaries
Search-R1. Recent advances, such as Search-R1, extend RAG to support multi-step reasoning with interleaved retrieval. In this paradigm, given a question $Q$, the LLM generates a reasoning trajectory $T=\{T_{1},T_{2},\ldots,T_{n}\}$. At each step $i$, the LLM (i) first generates a reasoning step $T_{i}$ based on the current information; (ii) then, it issues a search query and retrieves relevant documents; and (iii) finally, it judges whether to move on to the next reasoning step $T_{i+1}$ or generate the final answer AA based on the current content. The process alternates between reasoning and search.
@@ 3.2 Bidirectional information quantification
    Step-to-answer distance $d_{\text{T-A}}(T_{i})$ quantifies how much the current step $T_{i}$ contributes toward the final answer AA, indicating its solution progress; and
@@ 3.3 Multi-objective optimization with RL
represent the rewards derived from the step-to-answer and step-to-question distances at step ii.
@@
$J_{\text{GRPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot|x;\mathcal{R})}\big[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}$ (8)
@@
$-\beta D_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}]\big],$